OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (-6, 4) and (2, 0).

OpenStudy (anonymous):

@abb0t

OpenStudy (misssunshinexxoxo):

y − 4 = −one half(x + 6)

OpenStudy (misssunshinexxoxo):

@TheDragon do you know point slope intercept form?

OpenStudy (anonymous):

The choices are y - 4 = -1/2 (x + 6) y - 4 = 2 (x + 6) y + 6 = -1/2 (x - 4) y + 6 = 2 (x - 4)

OpenStudy (anonymous):

thanks, and kind of @misssunshinexxoxo

OpenStudy (abb0t):

For this one, you do \(\sf \large \color{blue}{\frac{y_2-y_1}{x_2-x_1}}\) and then plug it into the point-slope equation: So you get \(\large \frac{0-4}{2+6 } = \frac{-4}{8} = \frac{-1}{2} \) so your slope is \(-\frac{1}{2}\) now pick any point, i will pick second one y - 0 = \(\frac{1}{2}\)(x-2) or more simplified y = -\(\frac{1}{2}\)(x-2) which is the same thing as \(\sf \color{red}{y=-\frac{1}{2}x + 1}\)

OpenStudy (anonymous):

Thank you @abb0t can you help me on another question?